# Entropy-based Exploration Conduction for Multi-step Reasoning

### v1

@@ Abstract
In large language model (LLM) reasoning, multi-step processes have proven effective for solving complex tasks. However, the depth of exploration can significantly affect the reasoning performance. Existing methods to automatically decide the depth often bring high cost and lack flexibility, and thus undermine the model’s reasoning accuracy. To address these issues, we propose Entropy-based Exploration Depth Conduction (Entro-duction), a novel method that dynamically adjusts the exploration depth during multi-step reasoning by monitoring LLM’s output entropy and variance entropy. We employ these two metrics to capture the model’s current uncertainty and the fluctuation of uncertainty across consecutive reasoning steps. Based on the observed changes, the LLM selects whether to deepen, expand or stop exploration according to the probability. In this way, we balance the reasoning accuracy and exploration effectiveness. Experimental results across four benchmark datasets demonstrate the efficacy of Entro-duction. We further conduct experiments and analysis on the components of Entro-duction to discuss their contributions to reasoning performance.
@@
Jinghan Zhang1, Xiting Wang2thanks: Corresponding Author., Fengran Mo3, Yeyang Zhou4, Wanfu Gao5, Kunpeng Liu1 1Portland State University, 2Renmin University of China, 3University of Montreal, 4Uber, 5Jilin University {jinghanz,kunpeng}@pdx.edu xitingwang@ruc.edu.cn fengran.mo@umontreal.ca yeyang.zhou@uber.com gaowf@jlu.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable reasoning capabilities across various domains Brown et al. (2020); Touvron et al. (2023); Patterson et al. (2022); Fu et al. (2022); Wei et al. (2022). Meanwhile, facing complex tasks, they encounter challenges in generating accurate and effective multi-step reasoning. Specifically, LLMs may exhibit over-reasoning or under-reasoning, which both indicate the depth of exploration for a given problem does not match expectations Ahn et al. (2024); Mirzadeh et al. (2024); Huang and Chang (2022); Fu et al. (2023); Ying et al. (2024). This mismatch could lead to several issues: (1) inaccurate, insufficient, or redundant reasoning outcomes; (2) unnecessary computation costs Yeo et al. (2025); Yang et al. (2024); Lightman et al. (2023). These issues exist because i) LLMs lack evaluation and regulatory mechanisms for their reasoning process; ii) there are significant variations in the reasoning process required for different tasks, and LLMs are unable to accurately judge and adjust the depth of exploration for a task solely based on their training knowledge.
@@
Existing methods for optimizing the exploration depth of multi-step reasoning in LLMs fall into two types: outcome-based optimization and process-based optimization. Outcome-based optimization aligns the LLM’s reasoning exploration with human expectation after it has generated complete reasoning and reached conclusions Jin et al. (2024); Liu et al. (2024a); Ton et al. (2024); Yu et al. (2024); Gong et al. (2024). These approaches rely on post-training of LLMs, which is costly and does not provide an immediate improvement on the current task. Further, due to the diversity of reasoning tasks, the enhancement gain is task-specific. In process-based optimization, the LLMs supervise and evaluate their own reasoning process through self-critique or by labeling reasoning steps to enhance outputs Ma et al. (2025); Yang et al. (2024); Pan et al. (2023). The advantages of process-based optimization are the immediacy of optimization and its low cost. However, it requires the LLMs to have high reasoning capabilities and a substantial knowledge base. Moreover, since the process is usually opaque, it is difficult for humans to effectively supervise and interpret these optimization processes. And the model’s own illusions, biases, and errors may be reinforced during this process Stechly et al. (2023, 2024b); Liang et al. (2024b); Song et al. (2025).
@@
Given the existing challenges, we want to develop a method that guides LLMs to automatically, concisely, and transparently determine the appropriate depth of exploration based on task information and the model’s reasoning state. We aim to enable the model to look ahead during reasoning, plan dynamically, and decide whether further exploration is necessary to achieve optimal reasoning outcomes. Our target involves enhancing multi-step reasoning performance and reducing unnecessary exploration.
@@
To tackle these challenges, we propose Entropy-based Exploration Depth Conduction (Entro-duction), a novel method to help LLM assess the adequacy of exploration during multi-step reasoning processes, thus enhancing reasoning outcomes. Inspired by Entropy Uncertainty Measurement, we employ entropy changes in the LLM’s reasoning process to evaluate the model’s uncertainty of reasoning, which reflects the LLM’s reasoning confidence, and accordingly switch exploration strategy. Specifically, we use entropy and variance entropy as rewards to update the LLM’s probability distribution for its next exploratory action, whether to deepen, stop, or expand exploration. This distribution then drives a behavior selection mechanism that enhances reasoning when exploration is insufficient and avoids redundant reasoning when it is sufficient.
@@
    We propose Entro-duction to help LLMs dynamically evaluate the adequacy of exploration based on LLMs’ reasoning uncertainty to enhance reasoning performance and avoid unnecessary exploration.
@@
    We further design an entropy-based exploration behavior selection mechanism, which refers to LLMs’ expectancy and confidence in reasoning.
@@ 2 Related Work
When responding to queries, LLMs typically provide direct outputs. For complex questions, direct outputs often fail to deliver correct answers because they may overlook deeper logical connections or contextual information Xia et al. (2024); Minaee et al. (2024); Wu et al. (2021). Multi-step reasoning involves instructing LLMs to decompose and progressively address problems, breaking down complex tasks into smaller, manageable units to significantly enhance reasoning capabilities Chu et al. (2023). The simplest structure of multi-step reasoning is the Chain of Thought (CoT) Wei et al. (2022); Wang and Zhou (2024), which links reasoning steps by connecting distinct thoughts into a linear, coherent sequence. Li et al. (2024); Jin and Lu (2024); Sprague et al. (2024).
To enable more comprehensive exploratory reasoning, some studies based on CoT have developed structured reasoning methods, such as self-consistent CoT (CoT-SC), Complex CoT, and Tree-of-Thought (ToT) Wang et al. (2022); Zhang et al. (2024b); Yao et al. (2024); Liu et al. (2024b); Mo and Xin (2024); Zhang et al. (2024a). These methods are called reasoning structures. They guide LLMs to do multi-directional exploration of problem solution spaces for superior reasoning solutions by capturing more complex and varied logical relationships Xia et al. (2024); Stechly et al. (2024a); Liang et al. (2024a). However, the breadth and depth of these reasoning structures highly depend on predefined settings and vary greatly across different tasks, limiting their generalizability and flexibility.
@@
The depth of reasoning structures refers to the number of layers or steps in the reasoning process, namely, the number of reasoning steps an LLM must undertake before reaching a final answer Plaat et al. (2024); Gomez (2023). For any given task, the optimal number of reasoning layers often correlates with the task’s complexity and the level of detail required Zhang et al. (2024b). Current methods in determining these layers or optimize reasoning structures automatically. These methods include using reinforcement learning algorithms to optimize the number of reasoning steps or dynamically adjusting the depth of exploration during the reasoning processJin et al. (2024); Liu et al. (2024a); Hoffmann et al. (2022).
@@ 3.1 Problem Formulation
Given a reasoning task $Q$ and an LLM $\mathcal{R}$ as the reasoner, the multi-step reasoning process is to generate a reasoning structure $\mathcal{S}$. Structure $\mathcal{S}$ comprises directed links connecting sentence-level reasoning nodes. A reasoning chain is a unidirectional sequence that begins at an initial reasoning node and concludes at a terminal node:
@@ 3.4 Behavior Selection Mechanism
We introduce a mapping $\Phi:\mathbb{R}^{2}\rightarrow\mathcal{A}$ that assigns to each state $(\Delta H_{j},\,\Delta\sigma_{H,j}^{2})$ a “best” action $a_{j}^{*}$:
@@ 5 Conclusion
We develop a framework with dynamic depth adjustment strategies for LLMs. If not precisely calibrated, it might lead to suboptimal reasoning performance.. This may limit the Entro-duction method’s ability to adaptively balance exploration and exploitation in real-time. Besides, the experiments are conducted on four benchmark datasets on one Llama model, which may not provide a comprehensive view of the Entro-duction’s generalization capability across LLMs with varying sizes and pre-training processes. Moreover, the Entro-duction is mainly evaluated on specific tasks and these tasks cannot fully reflect the complexities of real-world scenarios where reasoning tasks can be variable and with more complex and external solution spaces. We left these potential explorations as future work.

### v2

@@ Abstract
Multi-step processes via large language models (LLMs) have proven effective for solving complex reasoning tasks. However, the depth of exploration of the reasoning procedure can significantly affect the task performance. Existing methods to automatically decide the depth often lead to high cost and a lack of flexibility. To address these issues, we propose Entropy-based Exploration Depth Conduction (Entro-duction), a novel method that dynamically adjusts the exploration depth during multi-step reasoning by monitoring LLM’s output entropy and variance entropy. We employ these two features to capture the model’s uncertainty of the current step and the fluctuation of uncertainty across consecutive reasoning steps. Based on the observed entropy changes, the LLM selects whether to deepen, expand, or stop exploration according to the probability, which facilitates the trade-off between the reasoning accuracy and exploration effectiveness. Experimental results across four benchmark datasets demonstrate the efficacy of Entro-duction.
@@
Jinghan Zhang1, Xiting Wang2thanks: Corresponding Author. Beijing Key Laboratory of Research on Large Models and Intelligent Governance. Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE., Fengran Mo3, Yeyang Zhou4, Wanfu Gao5, Kunpeng Liu1 1Portland State University, 2Gaoling School of Artificial Intelligence Renmin University of China Beijing, China, 3University of Montreal, 4Uber, 5Jilin University {jinghanz,kunpeng}@pdx.edu xitingwang@ruc.edu.cn fengran.mo@umontreal.ca yeyang.zhou@uber.com, gaowf@jlu.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable reasoning capabilities across various domains Brown et al. (2020); Touvron et al. (2023); Patterson et al. (2022); Fu et al. (2022); Wei et al. (2022). However, they would still encounter challenges in generating accurate and effective multi-step reasoning in terms of complex downstream tasks. Specifically, LLMs may exhibit over-reasoning or under-reasoning, which both imply the depth of exploration for a given problem does not match expectations Ahn et al. (2024); Mirzadeh et al. (2024); Huang and Chang (2022); Fu et al. (2023). This mismatch issue of reasoning path could lead to several issues: (1) inaccurate, insufficient, or redundant reasoning outcomes; (2) unnecessary computation costs Yeo et al. (2025); Yang et al. (2024); Lightman et al. (2023). These issues might be attributed to two aspects: i) the lack of evaluation and regulatory mechanisms for LLMs’ reasoning process; ii) there are significant variations in the reasoning process required for different tasks, and LLMs might not be unable to accurately judge and adjust the depth of exploration for a task solely based on their parametric knowledge during pre-training.
@@
Existing methods for optimizing the exploration depth of multi-step reasoning in LLMs can be categorized into two types: outcome-based optimization and process-based optimization. Outcome-based optimization aligns the LLMs’ reasoning exploration with human expectation after generating a complete reasoning path and approaching the final conclusions Jin et al. (2024); Liu et al. (2024a); Ton et al. (2024); Yu et al. (2024). These approaches rely on post-training techniques of LLMs, which are resource-intensive and do not provide an immediate improvement on the current task. Further, due to the diversity of reasoning tasks, the enhancement gain is task-specific. In process-based optimization, the LLMs supervise and evaluate their reasoning process through self-critique or by labeling reasoning steps to enhance outputs Ma et al. (2025); Yang et al. (2024); Pan et al. (2023). The advantages of the process-based optimization methods are their immediacy and low cost. However, the LLMs should have high reasoning capabilities and a substantial knowledge base. Moreover, since the process is usually opaque, it is difficult for humans to effectively provide supervision signals and interpret the optimization processes. The model’s illusions, biases, and errors may be reinforced during this process Stechly et al. (2023, 2024b); Liang et al. (2024b); Song et al. (2025).
@@
Given the existing challenges, we aim to develop a method that guides LLMs to automatically, concisely, and transparently determine the appropriate depth of exploration based on task information and the model’s reasoning state. The goal is to enable the model to look ahead during reasoning, plan dynamically, and decide whether further exploration is necessary to achieve optimal reasoning outcomes. The whole procedure involves enhancing multi-step reasoning performance and reducing unnecessary exploration.
@@
To tackle these challenges, we propose Entropy-based Exploration Depth Conduction (Entro-duction), a novel method to help LLMs assess the adequacy of exploration during multi-step reasoning processes, thus enhancing the outcomes of reasoning. Inspired by Entropy Uncertainty Measurement Coles et al. (2017); Farquhar et al. (2024); Zhang et al. (2024a); Rosenfeld et al. (1996), we employ entropy changes in the LLM’s reasoning process to evaluate its uncertainty of reasoning, which reflects the reasoning confidence, and accordingly switch exploration strategy. Specifically, we use entropy and variance entropy as rewards to update the LLM’s probability distribution for its next exploratory action, whether to deepen, stop, or expand exploration. This distribution subsequently guides a behavior selection mechanism that promotes reasoning when exploration is insufficient and avoids redundant reasoning when it is adequate.
@@
    We propose Entro-duction to help LLMs dynamically evaluate the adequacy of exploration based on their reasoning uncertainty to enhance reasoning performance and avoid unnecessary exploration.
@@
    We further design an entropy-based exploration behavior selection mechanism, which refers to LLMs’ expectancy and confidence in the reasoning procedure.
@@ 2 Related Work
When responding to queries, LLMs typically provide direct outputs. For complex questions, direct outputs often fail to deliver correct answers because they may overlook deeper logical connections or contextual information Xia et al. (2024); Minaee et al. (2024). Multi-step reasoning involves instructing LLMs to decompose and progressively address problems, breaking down complex tasks into smaller, manageable units to significantly enhance reasoning capabilities Chu et al. (2023). The simplest structure of multi-step reasoning is the Chain of Thought (CoT) Wei et al. (2022); Wang and Zhou (2024), which links reasoning steps by connecting distinct thoughts into a linear, coherent sequence Li et al. (2024); Jin and Lu (2024); Sprague et al. (2024).
To enable more comprehensive exploratory reasoning, some studies based on CoT have developed structured reasoning methods, such as self-consistent CoT (CoT-SC), Complex CoT, and Tree-of-Thought (ToT) Wang et al. (2022); Zhang et al. (2024c); Yao et al. (2024); Mo et al. (2024); Liu et al. (2024b); Mo and Xin (2024); Zhang et al. (2024b). These methods are called reasoning structures. They guide LLMs to do multi-directional exploration of problem solution spaces for superior reasoning solutions by capturing more complex and varied logical relationships Xia et al. (2024); Stechly et al. (2024a); Mo et al. (2023); Liang et al. (2024a). However, the breadth and depth of these reasoning structures highly depend on predefined settings and vary greatly across different tasks, limiting their generalizability and flexibility.
@@
The depth of reasoning structures refers to the number of layers or steps in the reasoning process, namely, the number of reasoning steps an LLM must undertake before reaching a final answer Plaat et al. (2024); Gomez (2023). For any given task, the optimal number of reasoning layers often correlates with the task’s complexity and the level of detail required Zhang et al. (2024c). Current methods in determining these layers or optimize reasoning structures automatically. These methods include using reinforcement learning algorithms to optimize the number of reasoning steps or dynamically adjusting the depth of exploration during the reasoning process Jin et al. (2024); Liu et al. (2024a); Hoffmann et al. (2022).
@@ 3.1 Problem Formulation
Given a reasoning task and an LLM $\mathcal{R}$ as the reasoner, the multi-step reasoning process is to generate a reasoning structure $\mathcal{S}$. Structure $\mathcal{S}$ comprises directed links connecting sentence-level reasoning nodes. A reasoning chain is a unidirectional sequence that begins at an initial reasoning node and concludes at a terminal node:
@@ 3.4 Behavior Selection Mechanism
We introduce a mapping function $\Phi:\mathbb{R}^{2}\rightarrow\mathcal{A}$ that assigns to each state $(\Delta H_{j},\,\Delta\sigma_{H,j}^{2})$ a “best” action $a_{j}^{*}$:
@@ 5 Conclusion
We develop a framework with dynamic depth adjustment strategies for LLMs. If not precisely calibrated, it might lead to suboptimal reasoning performance. This may limit the Entro-duction method’s ability to adaptively balance exploration and exploitation in real-time. Besides, the experiments are conducted on four benchmark datasets on one Llama model, which may not provide a comprehensive view of the Entro-duction’s generalization capability across LLMs with varying sizes and pre-training processes. Moreover, the Entro-duction is mainly evaluated on specific tasks and these tasks cannot fully reflect the complexities of real-world scenarios where reasoning tasks can be variable and with more complex and external solution spaces. We left these potential explorations as future work.
Acknowledgements
Dr. Xiting Wang is supported by the National Natural Science Foundation of China (NSFC) (NO. 62476279, NO. 92470205), Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China No. 24XNKJ18. Dr. Xiting Wang is supported by fund for building world-class universities (disciplines) of Renmin University of China and Public Computing Cloud, Renmin University of China.
Appendix